# Quality Time: Carbon-Aware Quality Adaptation for Energy-Intensive Services

### v2

@@ 4. Online Approach
For assessing the practicability of carbon-aware QoR adaptation under realistic conditions, we propose an online approach, which iteratively determines the optimal $A^{i}$ and $D^{i}$ for every interval $\mathcal{T}_{i}$. Its performance is limited by two factors:
@@
    the error of request and carbon intensity forecasts
@@
    the error introduced by approximating the NP-hard optimization problem
In our experiments, conducted with realistic forecasts and strict time limits for solving the MILP, we achieve emission savings of $82\pm 6$ relative to the upper-bound potential (Section 5.4).
4.1. Multi-Horizon Optimization
Based on the optimization problem introduced in Section 3.5, we propose a forecast-based multi-horizon optimization which determines the optimal deployment and load balancing for every $\mathcal{T}_{i}$.
Forecasting.  There exist various approaches for long-term load forecasting, such as Prophet (Taylor and Letham, 2017), DeepAR (Flunkert et al., 2017), or TBATS (De Livera et al., 2011). As the electricity mix of regions does not change abruptly, long-term forecasts for ACI can be usually estimated reasonably well by fitting daily, weekly, and annual seasonalities. However, especially in regions that are heavily dependent on weather-dependent renewable power production, additionally using short-term forecasts can provide significantly higher accuracy (Maji et al., 2023; Li et al., 2024c). For example, CarbonCast (Maji et al., 2023) reports mean absolute percentage errors (MAPE) of only 3-10 % for 24-hour forecasts.
Algorithm.  The approach is described in Algorithm 1. As the optimization problem is computationally hard, we approximate near-optimal solutions by either imposing time limits or terminating the optimization process once it reaches a solution within an acceptable proximity to optimality. To reduce the negative impact of these approximations on our decisions, we perform the optimization in two distinct steps with different levels of precision:
@@ 4.1. Multi-Horizon Optimization
  1. (1)
    A long-term optimization, executed every $\tau$ intervals (for example, every 24 hours), solves the MILP for the remainder of the year. Any $D^{i}$ and $A^{i}$ in the past are fixed, i.e. we only optimize $(D^{i},A^{i})_{i=\alpha}^{I-1}$, where $\alpha$ is the current interval. In this step, longer execution times and less optimal solutions are acceptable.
  2. (2)
    A short-term optimization, executed every interval, solves the problem within a fixed horizon [^2] We found that the validity period length $\gamma$ is a good indicator for horizon length, although it can be shorter or longer (as future values of $A^{i}$ are fixed by the long-term optimization). In Line 7 we assume $\gamma$ as the horizon.. As we only optimize over $(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma}$, this optimization involves significantly fewer variables and is more likely to find an optimal or near-optimal solution quickly. If no solution is found, we select an $A^{i}$ such that $\text{QoR}(\alpha,\omega)=1$, and determine the minimal $D^{i}$ that satisfies the Sufficient resources (12) constraint.
4.2. Automatic QoR Adaptation
The optimization problem introduced in Section 3.5 minimizes emissions under a given $\text{QoR}_{\text{target}}$. However, especially for best-effort users, another relevant question is: What are the expected annual emissions from this user group, and can we keep them within a predefined budget?
In Section 5.4, we show how running multiple long-term optimizations can provide service providers with valuable insights into expected annual emissions, allowing for adjustments to $\text{QoR}_{\text{target}}$ throughout the year. Building on this, we propose a variation of the online approach introduced in Section 4, which finds an optimal $\text{QoR}_{\text{target}}$ for best-effort users to stay within a carbon budget $B$:
(15) $\max\quad$ $\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=\alpha}^{\omega}$
@@ 4.2. Automatic QoR Adaptation
(16) $\sum_{i=\alpha}^{\omega}E^{i}\leq B$
In the multi-horizon optimization (Algorithm 1), we replace the long-term optimization in Line 5 with this objective:
(17) $(D^{i},A^{i})_{i=\alpha}^{I-1}\leftarrow\max_{(D^{i},A^{i})_{i=\alpha}^{I-1}}\left(\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=0}^{I-1}\right)$
Note, that we solve the optimization over $i=0,\dots,I-1$ (not just $i\geq\alpha$) as we must aggregate emissions over all intervals in (16). This has no performance implication, as we still only optimize the variables $(D^{i},A^{i})_{i=\alpha}^{I-1}$. Past values in $D^{i}$ and $A^{i}$ are fixed.
5. Experiments
We assess the potential of carbon-aware QoR adaptation by simulating the execution of an LLM inference service over the year 2023 using 8 real and synthetic request traces in 10 different regions.
We use Gurobi (Gurobi Optimization, LLC, 2024) to solve optimization problems. All simulations were conducted under the same conditions on an HPC cluster, with 16 cores allocated to each job.
Table 2. Upper-bound potential and carbon savings under realistic conditions in % for all request traces and regions at $\gamma=1$ week and $\text{QoR}_{\text{target}}=0.5$. Note, that other $\text{QoR}_{\text{target}}$ can lead to even better results (Figure X).
Region Artificial Real Synthetic Mean
Static Random Wiki (en) Wiki (de) Taxi Cell B Cell D Cell F
NL 8.2 / 7.4 8.5 / 6.3 7.6 / 6.5 10.2 / 9.3 11.3 / 8.7 7.8 / 6.7 7.6 / 6.6 8.0 / 7.0 8.7±1.3 / 7.3±1.0
CISO 7.0 / 6.8 7.5 / 6.0 7.5 / 7.0 10.0 / 9.4 11.3 / 9.2 6.6 / 6.1 6.3 / 5.9 6.8 / 6.4 7.9±1.7 / 7.1±1.3
ES 7.1 / 6.3 7.3 / 5.2 6.5 / 5.6 9.4 / 8.4 10.4 / 8.0 6.6 / 5.6 6.6 / 5.7 6.9 / 5.9 7.6±1.4 / 6.3±1.1
AU-QLD 7.9 / 7.8 8.4 / 7.0 7.1 / 6.5 6.1 / 5.8 8.8 / 6.9 7.6 / 7.0 6.9 / 6.4 8.0 / 7.5 7.6±0.8 / 6.9±0.6
DE 6.3 / 5.4 6.7 / 4.5 5.7 / 4.6 8.7 / 7.6 9.7 / 6.9 5.9 / 4.8 5.8 / 4.8 6.1 / 5.1 6.9±1.4 / 5.5±1.1
PL 4.1 / 3.8 4.5 / 2.9 3.5 / 2.9 6.7 / 6.2 7.5 / 5.7 3.8 / 3.3 3.7 / 3.2 4.1 / 3.5 4.7±1.4 / 3.9±1.2
ERCOT 4.2 / 3.1 4.6 / 2.6 3.6 / 2.4 6.3 / 5.1 7.4 / 5.8 3.9 / 2.8 3.7 / 2.5 4.1 / 3.0 4.7±1.3 / 3.4±1.2
SE 2.7 / 2.4 3.0 / 1.6 2.0 / 1.5 4.9 / 4.5 6.1 / 5.4 2.3 / 1.9 2.2 / 1.7 2.6 / 2.2 3.2±1.4 / 2.6±1.4
NYISO 2.3 / 1.8 2.7 / 1.1 1.7 / 1.1 4.6 / 4.3 5.6 / 4.6 1.9 / 1.5 1.8 / 1.3 2.2 / 1.7 2.8±1.4 / 2.2±1.3
PJM 1.8 / 1.7 2.3 / 0.8 1.3 / 1.0 4.1 / 4.0 5.5 / 4.3 1.6 / 1.4 1.4 / 1.1 1.8 / 1.6 2.5±1.4 / 2.0±1.3
5.1. Experimental Setup
As ACI is usually reported hourly, we define $\Delta=1$ hour.
[…]
Request traces.  For our analysis, we require 4 years of hourly request traces: 2020-2022 for fitting forecasting models and 2023 for the analysis itself. As such data are not publicly available for LLM inference servies, we show the effectiveness of our method across a diverse set of other real and synthetic request traces:
@@ 5.1. Experimental Setup
  • Real: Two real-life traces representing the English and German Wikipedia Pageview statistics (Wikimedia Foundation, [n. d.]), Wiki (en) and Wiki (de), and a trace based on the hourly taxi trips in New York City (Taxi and Commission, 2025), called Taxi.
  • Synthetic: Three generated traces, called Cell B, D, and F, which are based on instance event traces of Borg cluster cells at Google (Google, 2019).
All request trace datasets and our forecast methodology are explained in detail in Appendix B.
Carbon intensity traces.  We evaluate the potential for ten globally distributed regions:
  • In Europe, we consider Germany (DE), Spain (ES), Netherlands (NL), Poland (PL), and Sweden (SE)
  • In the US, we consider California (CISO), Texas (ERCOT), New York (NYISO), and the Pennsylvania-Jersey-Maryland Interconnection (PJM)
  • In Australia, we consider Queensland (AU-QLD)
All carbon intensity data was provided by ElectricityMaps (Maps, 2024). We normalized all time zones to align the time of day with the request traces. The temporal variability of regions is depicted in Figure 1. Our forecast methodology is explained in Appendix C.
@@ 5.2. Absolute Emissions
We first analyze the absolute annual emissions without carbon-aware QoR adaption, i.e. for $\gamma=1$ hour. The results reflect emission savings driven purely by increased energy demand from serving more requests on Llama 3.1 70B ($\text{QoR}_{\text{target}}\to 1$). Figure 4 shows the annual emissions for Wiki (en)—the trend is equivalent for all datasets. We observe very large differences across different regions: For example, Sweden’s energy mix relies almost entirely on renewable energy and nuclear power, making it one of the least carbon-intensive regions globally. Serving all requests with the 70B model results in just 245 tCO2, almost 27 times less than in Poland (4681 tCO2), where coal is the dominant energy source. As expected, the resulting emissions scale linearly with $\text{QoR}_{\text{target}}$.
Figure 4. Annual operational carbon emissions for Wiki (en) at different $\text{QoR}_{\text{target}}$ without carbon-aware QoR adaption.
Figure 5. Additional carbon savings resulting from increased validity periods, relative (top) and absolute (bottom) for $\text{QoR}_{\text{target}}=0.5$. Bigger validity periods yield larger gains.
@@ 5.3. Upper-Bound Potential
Carbon savings beyond energy efficiency.  In blue (left), Table 2 presents the relative additional carbon savings for $\text{QoR}_{\text{target}}=0.5$ at $\gamma=1$ week compared to $\gamma=1$ hour, where emissions directly depend on energy demand. We observe that through carbon-aware QoR adaptation, emissions drop by around 8 % in many regions. While request trace patterns exhibit little difference in potential, traces with a low volume of requests (Wiki (de) and Taxi) show increased potential, due to the higher impact of discrete scheduling decisions. For example, the volume of requests in Wiki (de) results in 10–25 active machines at a time, while Wiki (en) requires 80–200 active machines. In experiments with even fewer machines we managed to gain relative savings of more than 40 %—however, we do not analyze such cases further, as we focus on high-impact scenarios with many machines.
@@
Impact of validity periods.  Figure 5 reports the relative and absolute savings for $\text{QoR}_{\text{target}}=0.5$ and different validity periods. Increasing $\gamma$ to 8 hours yields only limited additional savings—less than 3 % across all settings. This is because average carbon intensity shows limited short-term variation, leaving little room for optimization. However, starting at $\gamma=24$ hours, we see a more notable potential of 5-8 % in some regions. Regional characteristics significantly influence the expected gains: For instance, while California’s carbon intensity primarily fluctuates on a daily basis due to its heavy reliance on solar energy, Germany exhibits more complex patterns with daily, weekly, and seasonal variations (Wiesner et al., 2021). As a result, extending $\gamma$ beyond 24 hours has minimal effect in California, while in Germany, larger $\gamma$ values lead to considerable improvements. In contrast, regions like Sweden, New York, and PJM exhibit very little temporal variation in carbon intensity.
Impact of premium users.  Figure 6 illustrates the loss of potential with a growing fraction of premium users at $\gamma=1$ week and $\text{QoR}_{\text{target}}=0.5$. We observe a slightly sub-linear decrease with a certain offset which depends on the overall request volume and, hence, the impact of discrete scheduling decisions. For reference, ChatGPT reached over 200 million active users in 2024 (Reuters, 2024), with fewer than 10 million paying customers (Tarantola, 2024). The resulting 5 % premium users decrease the potential savings by 5.6 % on average.
Figure 6. Impact of premium users on the potential savings.
Impact of $\text{QoR}_{\text{target}}$.  Figure 7 shows the potential for relative savings across different $\text{QoR}_{\text{target}}$. At $\text{QoR}_{\text{target}}=0$, we serve all requests using LLaMA 3.1 8B, leaving no flexibility for additional savings. Likewise, when $\text{QoR}_{\text{target}}$ is 1, all requests must be served by LLaMA 3.1 70B. A $\text{QoR}_{\text{target}}$ around 0.5 offers the greatest flexibility for carbon-aware QoR adaptation, resulting in the highest potential for carbon savings.
Periods of low service quality.  While increasing $\gamma$ enhances potential for carbon savings, it also increases the probability of prolonged periods with low QoR, illustrated in Figure 8. For example, when optimizing for $\text{QoR}_{\text{target}}=0.5$ at $\gamma=1$ week, no 1-week interval has a QoR below 0.5 (right). However, 10 % of all daily intervals exhibit a $\text{QoR}_{\text{target}}<0.48$ (left). We observe that for $\gamma=3$ months, users can experience multiple consecutive days of low QoR. Consequently, very large validity periods of more than a month are not practical in real-world scenarios.
5.4. Evaluation under Realistic Conditions
@@
We evaluated the practicability of carbon-aware QoR adaptation using the proposed multi-horizon optimization under real forecasts (see Appendix B as well as C) and approximated MILP solutions: Long-term optimizations are stopped at a $\text{MILP}_{\text{gap}}<0.1\,\%$ or after 30s; short-term optimizations at a $\text{MILP}_{\text{gap}}<0.1\,\%$ or after 10s.
[…]
Figure 9. Additional carbon savings under realistic conditions for Wiki (de) across all regions and $\text{QoR}_{\text{target}}=0.5$.
Outlook of annual emissions.  Long-term optimizations are computationally cheap when accepting near-optimal solutions—within just 30 seconds, the mean $\text{MILP}_{\text{gap}}$ reached 0.53 % accross all experiments. This allows us to generate a broad spectrum of solutions for varying $\text{QoR}_{\text{target}}$ and request forecasts, to represent, for instance, worst-case or best-case scenarios. Figure 10 demonstrates emission outlooks for Wiki (en) traces in California. For simulating best-case and worst-case request forecasts, we use the lower and upper bound of the 80 % prediction intervals provided by Prophet, respectively. Each week, we solve 33 long-term optimizations, covering the three forecasts and $\text{QoR}_{\text{target}}=\{0,0.1,...,1\}$.
(a) Annual emission outlook at three discrete points in time.
(b) Annual emission outlook; weekly for $\text{QoR}_{\text{target}}=0.5$.
Figure 10. Projected emissions (blue) based on long-term forecasts. The interval shows the range between best- and worst-case forecasts. Red indicates emissions already incurred.
Figure 10(a) shows the resulting outlooks a the start of the year, after 10 weeks, and after 40 weeks. The outlooks provide a comprehensive picture of the future, based on the currently available information. From the initial outlook (left), service providers may conclude that setting $\text{QoR}_{\text{target}}=0.5$ is a reasonable target to stay within an annual carbon budget of 1000 tCO2. Consequently, we optimize for $\text{QoR}_{\text{target}}=0.5$ for the remainder of the year.
Figure 10(b) shows that the initial estimate was overstated, although still in the range of the lower bound It depicts the weekly estimates for annual emissions when aiming for $\text{QoR}_{\text{target}}=0.5$. The upper and lower bound, defined by the worst-case and best-case forecasts, narrow over time as incurred emissions accumulate.
5.5. Automatic QoR Adaptation
Figure 11. Daily QoR for California (CISO), Texas (ERCOT), and Poland (PL). Our approach provides the most consistent QoR throughout the year.
Lastly, we illustrate the automatic adaptation of $\text{QoR}_{\text{target}}$ introduced in Section 4.2 on Wiki (en) in three exemplary regions. For each region, we define an annual budget based on the resulting emissions of the corresponding upper-bound experiment (Section 5.3) for $\gamma=24$ hours and $\text{QoR}_{\text{target}}=0.5$. Consequently, under perfect forecasts, the optimal approach would be to choose $\text{QoR}_{\text{target}}=0.5$ for every time step. We compare our approach to two greedy baselines that have access to the same forecasts as our approach:
  1. (1)
    Greedy (constant) divides the annual budget into equally sized hourly budgets for which it solves (15). Any remaining or overspent budget is evenly distributed across the remaining time intervals.
  2. (2)
    Greedy (weighted) also solves (15), but weights the hourly budgets based on forecasted load and carbon intensity. Every 24 hours, as request and carbon intensity forecasts are updated, weightings get adjusted.
Figure 11 shows the daily QoR of the greedy baselines and our automatic QoR adaptation. Since Greedy (constant) does not account for fluctuations in requests or carbon intensity, the provided QoR varies significantly throughout the year. Greedy (weighted) delivers more consistent QoR in all experiments but tends to fall back to either $\text{QoR}=1$ (NL) or $\text{QoR}=0$ (CISO and NYISO) toward the end of the year due to forecast errors. In the CISO and NYISO regions, it even exceeds the budget by 0.8 and 1.2 %, respectively. Our approach provides the most consistent service across all experiments as it periodically adjusts $\text{QoR}_{\text{target}}$ based on long-term optimizations. The annual budget is utilized to 98.2–99.8 %.
6. Discussion
This section highlights the impact of our results on energy-intensive services and discusses the limitations of this study.
Impact.  We emphasize the significant potential for carbon efficiency improvements in cloud services, based on our analysis in Section 5.2. ChatGPT, one of the most energy-intensive cloud services today, reached over 200 million active users in 2024 (Reuters, 2024). Assuming just 5 requests per user per day, this amounts to nearly 42 million requests per hour on average, which is an order of magnitude more than what we consider in our larger scenarios. Assuming performance similar to Llama 3.1 70B, we estimate that the annual operational carbon emissions of large-scale LLM service providers currently exceed multiple 10,000 tCO2—a figure set to rise with multi-modal AI models.
Furthermore, this study assumes ACI as the most likely metric to be implemented in official policy. However, and increasing part of the community is advocating to base load shifting decisions based on marginal signals (WattTime, 2022; Buchanan et al., 2023; Dodge et al., 2022), which—if implemented properly—better reflects the underlying complexities of electric grids. Yet, to this day, existing methodologies for MCI are based on predicting the marginal generator, which is ambiguous and error prone. Moreover, current reporting lacks spatial granularity to capture effects like energy curtailments from local grid congestion. We predict that with improved methodologies and more granular MCI reporting, carbon-aware QoR adaptation could have much greater potential, as MCI shows higher temporal variability than ACI, enabling the use of shorter validity periods. Our approach is directly applicable to MCI signals or any other scalar metric without requiring adaptations.
Limitations and future work.  This paper presents the first analysis on the potential of carbon-aware QoR adaptation in services, focusing on an LLM inference use case. While our approach is already applicable to more complex scenarios with multiple service quality tiers and heterogeneous infrastructure, such cases have not yet been investigated experimentally.
Second, our current modeling only considers the QoR for user groups, rather than individual users. This assumes a uniform distribution of users within each group and may lead to fairness concerns if the distribution is uneven. For example, ACI tends to peak in the evenings in most regions, meaning that users who usually access the service during these hours can experience a lower QoR than other users in their group. We consider fairness aspects with more granular user modeling as an area for future work.
Third, future work should explore the impact of adaptive QoR adjustments on user quality of experience and the resulting implications on user behavior. For instance, in the context of generative AI, reduced QoR might incentivize users to resubmit requests, potentially offsetting carbon savings. Unfortunately, to the best of our knowledge, there are no established behavioral models for generative AI usage. Developing such models would be valuable to better quantify the real-life impacts of sustainability approaches, including rebound effects.
7. Related Work
We survey related works on carbon-aware computing as well as carbon efficiency in energy-intensive cloud services.
Carbon-aware cloud services.  The large majority of approaches in carbon-aware computing focuses on shifting flexible batch processing towards times (Wiesner et al., 2021; Radovanovic et al., 2022; Fridgen et al., 2021; Lechowicz et al., 2023) and locations (Zheng et al., 2020; Lindberg et al., 2022; Wiesner et al., 2024) with clean energy. Hence, for batch jobs, the limitations and trade-offs of carbon-aware spatio-temporal workload shifting have been investigated thoroughly (Hanafy et al., 2024b; Sukprasert et al., 2024b; Hanafy et al., 2024a; Lechowicz et al., 2024). However, carbon efficiency in continuously running services that usually have little to no delay-tolerance has received only little attention so far.
Existing approaches for carbon-aware services primarily focus on load balancing across geo-distributed data centers and the respective capacity right-sizing under SLA constraints (Zhou et al., 2013, 2016). For example, Casper (Souza et al., 2023) is a carbon-aware scheduling and provisioning system for web services that minimizes the carbon footprint under latency constraints. CDN-Shifter (Murillo et al., 2024) explores carbon-aware spatial load shifting for content delivery networks. Caribou (Gsteiger et al., 2024) offloads serverless workflows across geo-distributed regions based on average carbon intensity. However, there are no existing approaches that address services which must operate in a single region due to data, security, or infrastructure constraints.
Sustainability in generative AI inference.  Although the carbon footprint of AI training has been an active topic of research for some time (Patterson et al., 2021; Dodge et al., 2022; Strubell et al., 2020; Dhar, 2020) only recently there has been an increasing focus on AI inference (Chien et al., 2023; Eilam et al., 2023; Samsi et al., 2023). Especially, the rapid adoption of LLMs has resulted in multiple studies that try to quantify their growing carbon footprint (Nguyen et al., 2024; Faiz et al., 2024; Luccioni et al., 2024). Many recent approaches target sustainability aspects of inference services throuh power-aware scheduling (Stojkovic et al., 2024, 2025; Qiu et al., 2025; Patel et al., 2024) or by considering their carbon and water footprint in geographic load balancing (Li et al., 2024d).
Service quality adaptation.  While service degradation is a common strategy to prevent system overload (Meza et al., 2023; Park et al., 2024), recent research has proposed designing applications with different service quality tiers as a means of improving sustainability (Vitali, 2022). In the context of carbon-aware LLMs, we identified only one approach, called Sprout (Li et al., 2024b), which optimizes the autoregressive generation process. Sprout returns shorter responses during periods of high carbon intensity, as the energy demand directly correlates with the number of generated tokens. Thereby, it effectively adapts QoR over time and can be further enhanced by leveraging the optimized infrastructure deployments and load balancing proposed in this paper.

### v3

@@ 3. Online Approach
A practical approach for carbon-aware QoR adaptation must cope with two sources of inaccuracy:
@@
    Approximation errors: Optimally solving the MILP over a large solution space is computationally infeasible. In our experiments, no annual-horizon instance solved within 1 hour, while the median runtime on daily-horizon instances was 1.2 seconds (see Appendix C).
@@
    Forecast errors: Long-term forecasts tend to be less accurate than short-term ones, particularly in regions with volatile renewable generation (Maji et al., 2023; Li et al., 2024b).
Based on these findings, we propose a multi-horizon optimization which periodically updates forecasts and determines an optimal (or close-to-optimal) deployment and load balancing for every interval $i$, described in Algorithm 1. We perform the optimization in two distinct steps:
  1. (1)
    A long-term optimization, ensures global feasibility, i.e., meeting QoR constraints. It is executed every $\tau$ intervals (e.g., every 24 hours) and solves the MILP for the remainder of the year. Any past $D^{i}$ and $A^{i}$ are fixed, so we only optimize $(D^{i},A^{i})_{i=\alpha}^{I-1}$, where $\alpha$ is the current interval. Solutions may be approximate due to solver time limits.
  2. (2)
    A short-term optimization adapts to recent forecast updates and improves suboptimal long-term decisions. It is executed every interval and solves the problem within a fixed horizon [^2] We found that the validity period length $\gamma$ is a good indicator for horizon length, although it can be shorter or longer (as future values of $A^{i}$ are fixed by the long-term optimization). In Line 7 we assume $\gamma$ as the horizon.. As we only optimize over $(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma}$, this optimization has a significantly smaller search space and is more likely to find an optimal solution quickly. If no solution is found, we select an $A^{i}$ such that $\text{QoR}(\alpha,\omega)=1$, and determine the minimal $D^{i}$ that satisfies Equation (5).
@@
4. Experiments
We evaluate our approach by simulating an LLM inference service over the year 2023 using 8 real and synthetic request traces in 10 different regions. All experiments were provided 16 cores of an HPC cluster, using Gurobi (Gurobi Optimization, LLC, 2024) for optimization.
@@ 4. Experiments
Request traces.  We require 4 years of hourly request traces: 3 for fitting forecasting models and 1 for the analysis itself. As such data are not publicly available for LLM inference services, we evaluate our method across a diverse set of other request traces:
@@
  • Human-Generated: Two real-world traces from unrelated domains, namely the English and German Wikipedia Pageview statistics (Wikimedia Foundation, [n. d.]), Wiki (en) and Wiki (de), and the hourly taxi trips in New York City (Taxi and Commission, 2025), called Taxi.
  • Synthetic: Three traces generated from models fitted to Google’s Borg cluster cells (Google, 2019), called Cell B, D, and F.
All request trace datasets and our forecast methodology are explained in detail in Appendix D.
Carbon intensity traces.  We evaluate on ten globally distributed regions: California, US (CISO), Texas, US (ERCOT), New York, US (NYISO), Germany (DE), Spain (ES), Netherlands (NL), Poland (PL), Sweden (SE), and Queensland, Australia (AU-QLD). All carbon intensity data was provided by ElectricityMaps (Maps, 2024). Our forecast methodology is explained in Appendix E.
@@ 4.1. Absolute Emissions
We first assess the absolute annual emissions without carbon-aware QoR adaption. Figure 3 shows the annual emissions for Wiki (de). As expected, emissions scale linearly with $\text{QoR}_{\text{target}}$, but we observe large differences across regions. For example, operating the service in Sweden emits almost 27 times less carbon than in Poland.
For reference, ChatGPT, one of the biggest LLM services today, reached over 200 million active users in 2024 (Reuters, 2024). Assuming just 5 requests per user per day, this amounts to nearly 42 million requests per hour—two orders of magnitude more than our Wiki (en) trace. Assuming similar performance as Llama 3.1 70B, we estimate that the annual operational carbon emissions of large-scale LLM service providers already exceed multiple 10,000 tCO2.
Figure 3. Annual operational carbon emissions for Wiki (de) at different $\text{QoR}_{\text{target}}$ without carbon-aware QoR adaption.
@@ 4.2. Upper-Bound Potential
Table 1 presents the potential for additional carbon savings for $\text{QoR}_{\text{target}}=0.5$ at $\gamma=1$ week. That is, instead of meeting the 50 % target hourly, we ensure that 50 % of requests are served by LLaMA 3.1 70B over each week. We observe that this increased flexibility enables emissions savings of around 8 % in many regions. While request trace patterns exhibit little difference in potential, low-volume traces (Wiki (de) and Taxi) offer greater savings due to the larger impact of discrete scheduling decisions. For example, the volume of requests in Wiki (de) results in 10–25 active machines at a time, while Wiki (en) requires 80–200 active machines. Note, that the potential depends heavily on $\text{QoR}_{\text{target}}$ as shown in Appendix F.
Table 1. Relative upper-bound carbon savings potential.
Region Artificial Human-Generated Synthetic Mean
Static Rnd. Wiki Wiki Taxi Cell Cell Cell
(en) (de) B D F
NL 8.2 8.5 7.6 10.2 11.3 7.8 7.6 8.0 8.7±1.3
CISO 7.0 7.5 7.5 10.0 11.3 6.6 6.3 6.8 7.9±1.7
ES 7.1 7.3 6.5 9.4 10.4 6.6 6.6 6.9 7.6±1.4
AU-QLD 7.9 8.4 7.1 6.1 8.8 7.6 6.9 8.0 7.6±0.8
DE 6.3 6.7 5.7 8.7 9.7 5.9 5.8 6.1 6.9±1.4
PL 4.1 4.5 3.5 6.7 7.5 3.8 3.7 4.1 4.7±1.4
ERCOT 4.2 4.6 3.6 6.3 7.4 3.9 3.7 4.1 4.7±1.3
SE 2.7 3.0 2.0 4.9 6.1 2.3 2.2 2.6 3.2±1.4
NYISO 2.3 2.7 1.7 4.6 5.6 1.9 1.8 2.2 2.8±1.4
PJM 1.8 2.3 1.3 4.1 5.5 1.6 1.4 1.8 2.5±1.4
Figure 4. Relative (top) and absolute (bottom) additional carbon savings for different validity periods of our proposed online approach (colored) compared to upper bound potential (gray). Larger validity periods enable higher savings.
@@
Figure 4 reports the relative and absolute saving potential for Wiki (de) at $\text{QoR}_{\text{target}}=0.5$ and different validity periods in gray. Increasing $\gamma$ to 8 hours yields only limited additional savings—less than 3 % across all settings. This is because average carbon intensity shows limited short-term variation, leaving little room for optimization. However, starting at $\gamma=24$ hours, we see a more notable potential of 5-8 % in some regions. Regional characteristics significantly influence the expected gains: For instance, while California’s carbon intensity primarily fluctuates on a daily basis due to its heavy reliance on solar energy, Germany exhibits more complex patterns with daily, weekly, and seasonal variations (Wiesner et al., 2021). As a result, extending $\gamma$ beyond 24 hours has minimal effect in California, while in Germany, larger $\gamma$ values lead to considerable improvements. In contrast, regions like Sweden, New York, and PJM exhibit little temporal variation in carbon intensity and have therefore a low savings potential of below 5 %.
4.3. Potential under Realistic Conditions
@@
We evaluate the practicability of carbon-aware QoR adaptation by applying our online approach under real forecasts (see Appendix D and E) and approximated MILP solutions: Long-term optimizations are stopped after 30s; short-term optimizations after 10s.
Figure 4 shows the gap to the upper-bound performance. Across all experiments, savings under realistic conditions reached $82\pm 6$ of the upper-bound potential. Notably, even in regions with highly unpredictable request patterns (Random and Cell B, D and F), the online optimization does not show a significant performance drop.